# Outward Sodium Current, Fine-structure Constant, and Ferroelectric Hysteresis Regimes in the Giant Squid Axon Propagating Action Potential: a Phase Space Approach

### v8

@@ Abstract
We derive a charge-conserving phase space cable equation for the propagating action potential, expressing ionic, capacitive and axoplasmic (membrane) currents as functions of membrane potential. Analysis of the ionic current during the recovery phase reveals an outward sodium current—unobserved in voltage-clamp experiments—arising from the axoplasmic modulated Ohmic current creating a positive charge gradient opposing the sodium concentration gradient.
@@ I Introduction
In sum, the axoplasmic current results in axoplasmic charge density gradients that resolve themselves in different ways along the action potential. At the leading edge of there is an accumulation of positive charges that keeps the propagation ongoing. The buildup of positive charge density just after the AP peak forces the outward sodium current against the sodium density gradient. The axoplasmic current maintains the coherence of the AP as it propagates.
@@
In the first part of our work, ANALYSIS - Part 1, we have rewritten the cable equation directly in Phase Space [34, p. 2]. The Phase Space cable equation yields several simple results. It shows that all three currents are proportional to $\Phi(V)=dV(t)/dt$ while the membrane current is independent of capacitance $C_{m}$. Since the capacitive current is proportional to the membrane’s capacitance, the ionic currents have contributions dependent and contributions not dependent on the capacitance. The phase Space cable equation also reveals the role of the propagation constant $k$ in the two crossings of the zero current axis by the total ionic current.
@@
We take it, as usual, that all ionic activation currents $J_{X}$, driven by ionic concentration density gradients, are described by three factors: maximum conductance $g_{X}$, reversal potential $V_{X}$ and fraction of open channels $Xo/X$. Phase space graphs of activation currents should reveal quasilinear segments and approximate values of reversal potentials, the corresponding maximum conductance (all channels open) and the time rates associated with the capacitive currents. We confirm the simple relationships between the propagation constant $k$, the maximum conductance $g_{X}$ and the corresponding time rates $\mu_{X}$ for the activation currents driven by the $X$ ionic density concentration gradients at the foot of the action potential and for the potassium current. Significantly, the graphs reveal an additional quasilinear segment of the outward current $J_{H}$ in the recovery region of the action potential. This current doesn’t have the simple relationships between the time rate, the maximum conductance and the propagation constant as those of the activation currents and we conclude that there is an additional driving force competing with the driving force of the sodium density gradient. In addition, the graph of the incoming sodium current doesn’t display an obvious reversal potential nor a quasilinear behavior associated with all channels open. Hodgkin and Huxley accounted for this inactivating behavior with the parameter $h$ that effectively decreases the fraction of open channels and thus competes with the driving force of sodium’s density gradient. We argue and demonstrate that the competing force causing the inactivation is the electrostatic charge gradient caused by the accumulation of positive charges by the axoplasmic modulated Ohm’s law current. We also conclude that the outward current $J_{H}$ is sodium’s deactivation (decreasing) current acting under a reversal potential result of the electrostatic potential overwhelming sodium’s concentration gradient. And, sodium H-channels close as the potential decreases in the recovery region. And notably, the maximum conductance of the H-channels is significantly smaller than the maximum conductance of M-channels as can be seen directly from ionic currents plots. This difference is the result of a structural change (continuous phase change) of the sodium channel at the peak of the action potential.
@@
The second part of our work, ANALYSIS Part 2, describes the fitting of all sodium, potassium activation, inactivation, deactivation and polarization ionic currents as products of three factors: maximum ionic channel conductance $g_{X}$, a combination of charge density gradients competing with ionic density gradients and a definition of the fraction of open channels given by the modified Avrami (mAvrami) equation, ${Xo(t)}{X}=1-e^{-\alpha_{X}[\mu_{X}(t-to_{X})]^{\theta_{X}}}$, where $\mu_{X}$ is the ionic time rate associated with capacitive currents and related to the maximum conductance and the propagation constant $k$, and where $\alpha_{X}$ are dimensionless, temperature independent, constants close to the value of the fine structure constant.
@@
Our fitting of experimental data predicts the existence of the outward sodium current during the recovery phase of the action potential and identifies continuous phase transitions at both the inception and peak of the action potential. These predictions should be regarded as hypotheses, to be tested through direct experimentation. Most intriguingly, and warranting further theoretical development, our fits suggest a possible role for the fine-structure constant as a universal scaling factor governing the possible tunneling of ions through channels. While the fitting itself does not constitute the proof of the fine-structure constant’s role in the passage of ions through channels, the fact that the fittings are consistently precise across different channels types and the full biological temperature range suggests that its role is, in some sense real. Consider the parallels with the story of Max Planck’s postulate regarding energy quanta.
@@
[…]
@@ II Analysis - Part 1
Figure 3: The experimental data Currents/$C_{m}\Phi(V)$ vs. V are presented as time trajectories. Note that the total ionic current $J_{I}$ is zero when $d\Phi(V)/dV$ is equal to the propagation constant $k$. Note that ionic current deals with specific ions traversing specific ionic channels while the membrane current deals with the net motion and distribution of Ohm’s charges in the axoplasm. There are four distinct segments of $d\Phi(V)/dV$, alternating positive and negative.  The first segment is positive. It anticipates and feeds the capacitive current and it is responsible for advancing the foot of the AP. This segment extends up to the inflection point of the AP rising edge (about 70 mV for Sweep 425) and it is mostly dedicated to the initial charging of the membrane’s capacitor at the foot of the AP allowing a small outward ionic current until $d\Phi(V)/dV=k$ when the outward ionic current crosses the zero current axis turning to incoming. In this segment the membrane current pushes axoplasmic positive charges in the direction of the AP motion down the AP rising edge slope.  The second segment is negative, it extends up the AP peak, and it corresponds to a decreasing rate of capacitive charge accumulation by the membrane current and it coincides with the first part of inactivation and the decrease of ionic conductance. It effectively flattens the potential across the membrane up to the AP peak.  The third segment is positive, it starts from null at the AP peak and grows very large as the action potential enters recovery and then decreases slowly practically to zero for a long stretch and then turns negative at about 30 mV. This behavior indicates further accumulation of positive charges in the axoplasm immediately after the AP peak leading to the potential at which the derivative $d\Phi(V)/dV=k$ and at which the ionic current turns from incoming to outgoing. The electrical gradient of positive charges across the membrane eventually overwhelms sodium ion’s concentration gradient by pushing positive ions outward, to be later shown to be sodium ions traversing through open sodium channels.  The fourth segment is negative and it extends up to the end of the action potential as the other outward ionic current (potassium current) decreases to zero by reaching its reversal potential. It corresponds to the membrane current retracting of the AP by decreasing the accumulation of axoplasmic positive charges to zero.
@@
Fig. 2 and Fig. 3 exhibit the complex behavior of the modulated axoplasmic axial Ohm’s current, the so-called membrane current. Plots of currents $k(C_{m}\Phi,J_{m},J_{I})$/$C_{m}\Phi$ vs. Potential (Fig. 3) show explicitly that all three currents are proportional to $\Phi(V)$ and that the ionic current crosses the zero current axis when $d\Phi(V)/dV=k$.
@@
    The quasilinear region straddling the first voltage point at which the ionic current crosses the zero current axis is very short and it is displayed in Fig. Fig. 7 for Sweep 425. The trend line of the capacitive current portion intercepts the zero axis current at a negative potential $V=-3mV$ and the trend lines of the corresponding ionic and membrane currents intercept at a positive current above $V=-3mV$. Trend lines are not shown in the SM Fig. 21. Similar situation involving the intercepts of the three trend lines will be shown to occur at the second quasilinear region straddling the potential at which total ionic current crosses the zero current axis (See Fig.  9 and SM Fig.  22).
@@ III Analysis - Part 2
where $to_{X}$ is the inception time, the time when the ionic channels start opening, and $A_{X}$ is the Avrami parameter and $\theta_{X}$ is the dimensionless Avrami exponent. The Avrami equation is best known for describing isothermal phase changing in solids and crystallization processes. The particular value of the exponent $\theta_{X}=4$ comes from from three dimensions of growth and one representing a constant nucleation rate. Avrami parameter $A_{X}$ is typically very temperature dependent. In particular, for incoming sodium’s current $A_{M}$ varies exponentially from about 60 at $1^{o}$ to 15000 at $25^{o}$.
Figure 5: Avrami parameters $A_{M}$, $A_{H}$ and $A_{N}$ vs. $({\mu_{M}})^{\theta_{M}}$, $({\mu_{H}})^{\theta_{H}}$ and $({\mu_{N}})^{\theta_{N}}$ respectively. (a) $A_{M}~{}vs.~{}(\mu_{M})^{\theta_{M}}$. Points correspond to sweeps at 11 different temperatures. The average value the parameter $\alpha_{M}$ is 0.0060, however the slope is $\alpha_{M}=0.0074$. While the time rate $\mu_{M}$ grows exponentially as function of temperature, $\theta_{M}\approx 4$ and $\alpha_{M}$ are practically temperature independent and inform about sodium M-channel structure. (b) $A_{H}~{}vs.~{}({\mu_{H}})^{\theta_{H}}$. Points correspond to Sweeps 170, 220, 327, 525, 630 and 695. The correlation $A_{H}\propto{\mu_{H}}^{\theta_{H}}$ also holds for the deactivation current $J_{H}$. Sodium’s H-channel time rate and conductance are different from the M-Sodium channel’s parameters as result of the continuous phase change from the M channel structure to the H structure. (c) $A_{N}~{}vs.~{}(\mu_{N})^{\theta_{N}}$. Similarly, for the N-channel the average value of the mAvrami parameter $\alpha_{N}$ is 0.01. While $\mu_{N}$ grows exponentially as function of temperature, $\alpha_{N}$ and $\theta_{N}\approx 3$ are practically temperature independent and inform about potassium N-channel structure. The values of the three slopes $\alpha_{M}$, $\alpha_{H}$ and $\alpha_{N}$ are very close to the value of he fine-structure constant $\alpha$. There is a small number of dimensionless and temperature independent parameters in physics. We conclude that all three $\alpha_{X}$ are equal to the FSC $\alpha$. In consequence, in what follows we have seeded the value of the FSC in all the fittings of the experimental data.
@@
Figure 6: Fittings of ionic currents with seeded values for fractions of open channels $\alpha=0.007297...$, $\theta_{M}=3.78$, $\theta_{H}=3$ and $\theta_{N}=3$. See SM Fig. 20 for the table of parameters. (a) $A_{M}~{}vs.~{}(\mu_{M})^{\theta_{M}}$. Points correspond to sweeps at 11 different temperatures. While the time rate $\mu_{M}$ grows exponentially as function of temperature, $\theta_{M}=3.78$ and $\alpha$ are temperature independent and inform about sodium M-channel structure. (b) $A_{H}~{}vs.~{}({\mu_{H}})^{\theta_{H}}$. Points correspond to Sweeps 170, 220, 327, 525, 630 and 695. The correlation $A_{H}=$$\alpha({\mu_{H}})^{\theta_{H}}$ also holds for the deactivation current $J_{H}$. H-Sodium channel time rate and conductance are different from M-Sodium channel’s parameters as result of the continuous phase change from M channel structure to the H structure. (c) $A_{N}~{}vs.~{}(\mu_{N})^{\theta_{N}}$. While $\mu_{N}$ grows exponentially as function of temperature, $\alpha$ and $\theta_{N}=3$ are temperature independent and inform about potassium N-channel structure.
@@
Also listed are the capacitive current $C_{m}\Phi_{K}$ corresponding to ionic current $J_{K}$ at the foot of the action potential, sodium and potassium capacitive currents $C_{m}\Phi_{M}$ and $C_{m}\Phi_{N}$ that are structured as a product of the time rate, the driving force and some function of the fraction of open channels. In particular for the M-channels this function is approximately the fraction of open channels exponentiated to 1/3, $(Mo/M)$1/3. The exponent 1/3 is an ad hoc parameter that yields almost the same mAvrami or Phase Space parameters for the fraction $Mo/M$ in Eq. 17c and Eq. 17d. Note that the region of fitting covers the potential from the inflection of the action potential to the start of the negative resistance of the action potential, allowing for a small inaccuracy of the capacitive activation current due to the ad hoc choice of the parameter (1/3).
@@
Figure 8: Following currents are shown: total experimental ionic and capacitive currents from Eq. (3) with their 6th order polynomial fits that avoid the polarization segments, sodium’s $J_{H}$ and potassium’s $J_{N}$ ionic mAvrami fits with their quasi linear segments matching the polynomial fits. Membrane current is shown as $J_{m}(V)$=$(1/k)C_{m}\Phi(V)d\Phi(V)/dV$ of the 6th order polynomial fit of $\Phi(V)$ plus a small experimental part close to the AP peak. Recovery polarization currents are not shown. The ionic polarization current, within few mV of the action potential peak, is displayed in Fig. 9. The linear slope of the $J_{H}(V)$ ionic current and the linear slope of the corresponding capacitive current $C_{m}\Phi_{H}(V)$ intercept the zero current axis at different points while the membrane doesn’t intercept the zero current axis in the vicinity of the AP peak. All three potassium’s activation currents intercept the zero current approximately at the same value as read from the graph. The two linear segments of the capacitive current correspond to rate constants of sodium H-channel $\mu_{H}$ and potassium N-channel $\mu_{N}$. The slope of $J_{H}$ linear segment is $-g_{H}$ where $g_{H}$ is the maximum conductance of the sodium H-channel. The slope of the ionic current $J_{N}$ is the maximum conductance $g_{N}$. $\bf{Note:}$ When fitting the PhS currents we did not require that $Vo_{N}$ and $Vo_{H}$ correspond to $to_{N}$ and $to_{H}$ respectively as we did for the rising edge $Vo_{M}$.
@@
The post inception ionic polarization current covers the region from the inception to the inflection of the AP. It starts outgoing and decreasing as it crosses the zero current axis and then turns incoming furthering the depolarization of the membrane and thus furthering the opening of sodium M-channels. The post inception ionic polarization current has a small maximum conductance $g_{P1}$ = $g_{P2}$. The reversal potential $V_{P1}$ = $V_{P2}$ decreases from about 40 mV to about 30 mV as temperature increases and it is much closer to the resting potential $V=0$ than to sodium reversal potentials $J_{H}$ $<$ $J_{M}$.
@@
Note that Eq. 17f is only approximately valid because the first and the fourth segments of $P{io}/Pi$ include the polarization flips at the inception point and at the peak of the action potential $V_{p}$ respectively (See Fig. 16 and SM Fig. 25). Similarly for Eq. 17g which includes the polarization flip at the peak of the action potential.
@@
The extension of the polynomial trend line of the quasilinear portion of the recovery capacitive current $C_{m}\Phi$approx $C_{m}\Phi_{H}$ intercepts the zero current axis at the point $V\approx V_{M}$ ) (See Fig. 9 and and Fig. 22). As we later apply the parsing of the rising edge ionic current into the activation mAvrami sodium current $J_{M}$ and the corresponding ionic polarization currents, our fits find that the reversal potential of the incoming activation current is $V_{M}$. The fact that reversal potentials of $J_{M}$, $J_{mM}$, $C_{m}\Phi_{M}$ and $C_{m}\Phi_{H}$ are the same supports the fact that the outgoing current $J_{H}$ is in fact sodium current. This result is also independent of any other assumption and it is solely determined by the charge conserving cable equation.
However, the current $J_{H}$ intercepts the zero current axis at $V=V_{H}$ that is significantly lower than the reversal potential $V_{M}$, and its polynomial trend line extension intercepts the slope of the recovery linear segment of the membrane current $J_{mH}$ at the potential $V_{M}$ but below the zero current axis. Fig. 10 displays the same behavior at all temperatures. These facts support the conclusion that $J_{H}$ is in fact sodium current with its reversal potential modified from $V_{M}$ to $V_{H}$ by the inactivation currents and the continuous phase change of the sodium M-channel symmetry into the sodium H-channel symmetry while keeping almost all the mAvrami channels open (see Fig. 15).
Linear segments of $J_{K}$ and $J_{N}$ and $J_{H}$, apparent at once in Figures 7, 8, 9, and SM Figures 21, 23, correspond to fractions of open channels equal to one; i.e. all channels open. The slopes of the capacitive linear segments correspond to time rate parameters $\mu_{K}$, $\mu_{M}$, $\mu_{H}$, $\mu_{N}$, and the slopes of the ionic current linear segments correspond to maximum conductance $g_{K}$, $g_{M}$, $g_{H}$ and $g_{N}$. Maximum conductance and time rates vs. Temperature are plotted in Fig. 12.
@@
Figure 11: The rising edge capacitive, membrane, ionic and corresponding polarizations, and mAvrami fitted currents are plotted against time. The three mAvrami fitted currents $C_{m}\Phi_{M}$, $J_{mM}$, $J_{M}$, and the $Mo(t)/M$ curve begin and intersect the zero-current axis at the inception point $t=to_{M}$, the time at which M-sodium channels start to open. Capacitive and membrane polarization currents $C_{m}\Phi_{MP}$ and $J_{mMP}$, start with the polarization flip at $t=to_{M}$ and then decay. The surface under the first segment of capacitive polarization curve $C_{m}\Phi_{MP}$ is equal to the polarization charge $Q_{P}\approx 10x10^{-9}$ Coulomb/cm2 moved across the membrane (a.k.a. gating charge). The flip from H-channel symmetry to M-channel symmetry symmetry and the transfer of charges across but within the membrane in conjunction with the ionic polarization current precedes the opening of sodium channels. In the present theory the motion of gating charges tapers off at approximately the maximum rate of rise of the action potential at about 60 mV when the fraction of open sodium channels is about $20\%$ The value of the charge $Q_{P}$ remains approximately the same at all temperatures. Towards the peak of the action potential, preceding the flip from M-channel symmetry to H-channel symmetry at the peak, there is an opposite charge transfer across and also within the membrane amounting to $q_{P}\approx-2.6x10^{-9}$ Coulomb/cm2.
@@
Linear segments of ionic currents $J_{K}$, $J_{M}$ and $J_{N}$ intersect the zero current axis at the same point as the corresponding linear segments of capacitive and membrane currents. The cable Eq. (3) yields for these linear segments the relationships between maximum conductance and corresponding time rate. See also the reference [45]:

### v9

# Physics of the Propagated Action Potential

@@ Abstract
We derive a charge-conserving phase space cable equation for the propagating action potential, expressing ionic, capacitive and axoplasmic (membrane) currents as functions of membrane potential. The new equation’s prediction of ionic current crossing the zero current axis twice as function of the propagation constant is confirmed by the Rosenthal-Bezanilla experimental data. Analysis of the ionic current during the recovery phase reveals an outward sodium current—unobserved in voltage-clamp experiments—arising from the axoplasmic modulated Ohmic current creating a positive charge gradient opposing the sodium concentration gradient.
@@ I Introduction
In sum, the axoplasmic current results in axoplasmic charge density gradients that resolve themselves in different ways along the action potential. At the leading edge of the action potential there is an accumulation of positive charges feeding the membrane’s capacitor that keeps the propagation ongoing. The buildup of positive charge density just after the AP peak forces the outward sodium current against the sodium density gradient. The axoplasmic current maintains the coherence of the AP as it propagates.
@@
In the first part of our work, ANALYSIS - Part 1, we have rewritten the cable equation directly in Phase Space [34, p. 2]. The Phase Space cable equation yields several simple results. It shows that its three currents and Ohm’s current are proportional to $\Phi(V)=dV(t)/dt$ while the membrane current is independent of capacitance $C_{m}$. Since the capacitive current is proportional to the membrane’s capacitance, the ionic currents have contributions dependent and contributions not dependent on the capacitance. The phase Space cable equation also reveals the role of the propagation constant $k$ in the two crossings of the zero current axis by the total ionic current.
@@
We take it, as usual, that all ionic activation currents $J_{X}$, driven by ionic concentration density gradients, are described by three factors: maximum conductance $g_{X}$, reversal potential $V_{X}$ and fraction of open channels $Xo/X$. Phase space graphs of activation currents should reveal quasilinear segments and approximate values of reversal potentials, the corresponding maximum conductance (all channels open) and the time rates associated with the capacitive currents. We confirm the simple relationships between the propagation constant $k$, the maximum conductance $g_{X}$ and the corresponding time rates $\mu_{X}$ for the activation currents driven by the $X$ ionic density concentration gradients at the foot of the action potential and for the potassium current. Significantly, the graphs reveal an additional quasilinear segment of the outward current $J_{H}$ in the recovery region of the action potential. This current doesn’t have the simple relationships between the time rate, the maximum conductance and the propagation constant as those of the activation currents and we conclude that there is an additional driving force competing with the driving force of the sodium density gradient. In addition, the graph of the incoming sodium current doesn’t display an obvious reversal potential nor a quasilinear behaviour associated with all channels open. Hodgkin and Huxley accounted for this inactivating behaviour with the parameter $h$ that effectively decreases the fraction of open channels and thus competes with the driving force of sodium’s density gradient. We argue and demonstrate that the competing force causing the inactivation is the electrostatic charge gradient caused by the accumulation of positive charges by the axoplasmic modulated Ohm’s law current (aka membrane current). Ohm’s current moves positive charges down potential and negative charge up potential. Since Ohm’s current is not constant along the action potential its derivative (the membrane current) is proportional to axoplasmic charge density change along he action potential that parses itself between the capacitive and ionic currents. We also conclude that the outward current $J_{H}$ is sodium’s deactivation (decreasing) current acting under a reversal potential result of the electrostatic potential overwhelming sodium’s concentration gradient. And, sodium H-channels close as the potential decreases in the recovery region. And notably, the maximum conductance of the H-channels is significantly smaller than the maximum conductance of M-channels as can be seen directly from ionic currents plots. This difference is the result of a structural change (continuous phase change) of the sodium channel at the peak of the action potential.
@@
The second part of our work, ANALYSIS Part 2, describes the fitting of all sodium, potassium activation, inactivation, deactivation and polarization ionic currents as products of three factors: maximum ionic channel conductance $g_{X}$, a combination of charge density gradients competing with ionic density gradients and a definition of the fraction of open channels given by the modified Avrami (mAvrami) equation, ${Xo(t)}{X}=1-e^{-\alpha_{X}[\mu_{X}(t-to_{X})]^{\theta_{X}}}$, where $\mu_{X}$ is the ionic time rate associated with capacitive currents, and where $\alpha_{X}$ are dimensionless, temperature independent, constants close to the value of the fine structure constant.
@@
Our fitting of experimental data predicts the existence of the outward sodium current during the recovery phase of the action potential and identifies continuous phase transitions at both the inception and peak of the action potential. These predictions should be regarded as hypotheses, to be tested through direct experimentation. Most intriguingly, and warranting further theoretical development, our fits suggest a possible role for the fine-structure constant as a universal scaling factor governing the possible tunnelling of ions through channels. While the fitting itself does not constitute the proof of the fine-structure constant’s role in the passage of ions through channels, the fact that the fittings are consistently precise across different channels types and the full biological temperature range suggests that its role is, in some sense real. Consider the parallels with the story of Max Planck’s postulate regarding energy quanta.
@@
We parse the total ionic, capacitive and membrane currents into ionic activation/deactivation currents described by mAvrami currents accounting for the bulk of the experimental data and the corresponding small differences between the total experimental currents and the mAvrami currents, called here, tinder and inactivation currents, respectively adding or opposing the mAvrami ionic current. In what follows, we call the tinder and inactivation currents collectively as polarization currents: $\Sigma_{i}$$J_{Pi}$. Ionic currents traversing channels are driven by gradients of ionic concentration across the membrane. Polarization currents are caused by the motion of dipole charges within the membrane opening and closing the channels; or caused by modulation of the charge concentration throughout the axoplasm by the membrane current that is proportional to the product of Ohm’s current and its derivative; or caused by structural continuous phase symmetry change of the channel. We fit them with mAvrami equations with effective reversal potentials and time rates and call these currents mAvrami polarization currents.
@@ II Analysis - Part 1
[…]
@@
Fig. 2 and Fig. 3 exhibit the complex behavior of the modulated axoplasmic axial Ohm’s current, the so-called membrane current. Plots of currents $k(C_{m}\Phi,J_{m},J_{I})$$C_{m}\Phi$ vs. Potential (Fig. 3) show explicitly that all three currents are proportional to $\Phi(V)$ and that the ionic current crosses the zero current axis when $d\Phi(V)/dV=k$.
@@
    The quasilinear region straddling the first voltage point at which the ionic current crosses the zero current axis is very short and it is displayed in Fig. Fig. 7 for Sweep 425. The trend line of the capacitive current portion intercepts the zero axis current at a negative potential V=3mVV=-3mV and the trend lines of the corresponding ionic and membrane currents intercept at a positive current above V=3mVV=-3mV. Trend lines are not shown in the SM Fig. 21. Similar situation involving the intercepts of the three trend lines will be shown to occur at the second quasilinear region straddling the potential at which total ionic current crosses the zero current axis (See Fig.  9 and SM Fig.  22).
@@ III Analysis - Part 2
where $to_{X}$ is the inception time, the time when the ionic channels start opening, and $A_{X}$ is the Avrami parameter and $\theta_{X}$ is the dimensionless Avrami exponent. The Avrami equation is best known for describing isothermal phase changing in solids and crystallization processes. The particular value of the exponent $\theta_{X}=4$ comes from from three dimensions of growth and one representing a constant nucleation rate. Avrami parameter $A_{X}$ is typically very temperature dependent. In particular, for incoming sodium’s current $A_{M}$ varies exponentially from about 60 at $1^{o}$C to 15000 at $25^{o}$C.
Figure 5: Avrami parameters $A_{M}$, $A_{H}$ and $A_{N}$ vs. $({\mu_{M}})^{\theta_{M}}$, $({\mu_{H}})^{\theta_{H}}$ and $({\mu_{N}})^{\theta_{N}}$ respectively. (a) $A_{M}~vs.~(\mu_{M})^{\theta_{M}}$. Points correspond to sweeps at $1$ different temperatures. The average value the parameter $\alpha_{M}$ is 0.0060, however the slope is $\alpha_{M}=0.0074$. While the time rate $\mu_{M}$ grows exponentially as function of temperature, $\theta_{M}\approx 4$ and $\alpha_{M}$ are practically temperature independent and inform about sodium M-channel structure. (b) $A_{H}~vs.~({\mu_{H}})^{\theta_{H}}$. Points correspond to Sweeps 170, 220, 327, 525, 630 and 695. The correlation $A_{H}\propto{\mu_{H}}^{\theta_{H}}$ also holds for the deactivation current $J_{H}$. Sodium’s H-channel time rate and conductance are different from the M-Sodium channel’s parameters as result of the continuous phase change from the M channel structure to the H structure. (c) $A_{N}~vs.~(\mu_{N})^{\theta_{N}}$. Similarly, for the N-channel the average value of the mAvrami parameter $\alpha_{N}$ is 0.01. While $\mu_{N}$ grows exponentially as function of temperature, $\alpha_{N}$ and $\theta_{N}\approx 3$ are practically temperature independent and inform about potassium N-channel structure. The values of the three slopes $\alpha_{M}$, $\alpha_{H}$ and $\alpha_{N}$ are very close to the value of he fine-structure constant $\alpha$. There is a small number of dimensionless and temperature independent parameters in physics. We conclude that all three $\alpha_{X}$ are equal to the FSC $\alpha$. In consequence, in what follows we have seeded the value of the FSC in all the fittings of the experimental data.
@@
Figure 6: Fittings of ionic currents with seeded values for fractions of open channels $\alpha=0.007297$..., $\theta_{M}=3.78$, $\theta_{H}=3$ and $\theta_{N}=3$. See SM Fig. 20 for the table of parameters. (a) $A_{M}~vs.~(\mu_{M})^{\theta_{M}}$. Points correspond to sweeps at $1$ different temperatures. While the time rate $\mu_{M}$ grows exponentially as function of temperature, $\theta_{M}=3.78$ and $\alpha$ are temperature independent and inform about sodium M-channel structure. (b) $A_{H}~vs.~({\mu_{H}})^{\theta_{H}}$. Points correspond to Sweeps 170, 220, 327, 525, 630 and 695. The correlation $A_{H}=$$\alpha({\mu_{H}})^{\theta_{H}}$ also holds for the deactivation current $J_{H}$. H-Sodium channel time rate and conductance are different from M-Sodium channel’s parameters as result of the continuous phase change from M channel structure to the H structure. (c) $A_{N}~vs.~(\mu_{N})^{\theta_{N}}$. While $\mu_{N}$ grows exponentially as function of temperature, $\alpha$ and $\theta_{N}=3$ are temperature independent and inform about potassium N-channel structure.
@@
Also listed are the capacitive current $C_{m}\Phi_{K}$ corresponding to ionic current $J_{K}$ at the foot of the action potential, sodium and potassium capacitive currents $C_{m}\Phi_{M}$ and $C_{m}\Phi_{N}$ that are structured as a product of the time rate, the driving force and some function of the fraction of open channels. In particular for the M-channels this function is approximately the fraction of open channels exponentiated to 1/3, (Mo/M)(Mo/M)1/3. The exponent 1/3 is an ad hoc parameter that yields almost the same mAvrami or Phase Space parameters for the fraction $Mo/M$ in Eq. 17c and Eq. 17d. Note that the region of fitting covers the potential from the inflection of the action potential to the start of the negative resistance of the action potential, allowing for a small inaccuracy of the capacitive activation current due to the ad hoc choice of the parameter (1/3).
@@
Figure 8: Following currents are shown: total experimental ionic and capacitive currents from Eq. (3) with their 6th order polynomial fits that avoid the polarization segments, sodium’s $J_{H}$ and potassium’s $J_{N}$ ionic mAvrami fits with their quasi linear segments matching the polynomial fits. Membrane current is shown as $J_{m}(V)$$(1/k)C_{m}\Phi(V)d\Phi(V)/dV$ of the 6th order polynomial fit of $\Phi(V)$ plus a small experimental part close to the AP peak. Recovery polarization currents are not shown. The ionic polarization current, within few mV of the action potential peak, is displayed in Fig. 9. The linear slope of the $J_{H}(V)$ ionic current and the linear slope of the corresponding capacitive current $C_{m}\Phi_{H}(V)$ intercept the zero current axis at different points while the membrane doesn’t intercept the zero current axis in the vicinity of the AP peak. All three potassium’s activation currents intercept the zero current approximately at the same value as read from the graph. The two quasilinear segments of the capacitive current correspond to rate constants of sodium H-channel $\mu_{H}$ and potassium N-channel $\mu_{N}$. The slope of $J_{H}$ linear segment is $g_{H}$ where $g_{H}$ is the maximum conductance of the sodium H-channel. The slope of the ionic current $J_{N}$ is the maximum conductance $g_{N}$. $\bf{Note:}$ When fitting the PhS currents we did not require that $Vo_{N}$ and $Vo_{H}$ correspond to $to_{N}$ and $to_{H}$ respectively as we did for the rising edge $Vo_{M}$.
@@
The post inception ionic polarization current covers the region from the inception to the inflection of the AP. It starts outgoing and decreasing as it crosses the zero current axis and then turns incoming furthering the depolarization of the membrane and thus furthering the opening of sodium M-channels. The post inception ionic polarization current has a small maximum conductance $g_{P1}$ = $g_{P2}$. The reversal potential $V_{P1}$ = $V_{P2}$ decreases from about 40 mV to about 30 mV as temperature increases and it is much closer to the resting potential $V=0$ than to sodium reversal potentials $J_{H}$ << $J_{M}$.
@@
Note that Eq. 17f is only approximately valid because the first and the fourth segments of Pio/PiP{io}/Pi include the polarization flips at the inception point and at the peak of the action potential $V_{p}$ respectively (See Fig. 16 and SM Fig. 25). Similarly for Eq. 17g which includes the polarization flip at the peak of the action potential.
@@
The extension of the polynomial trend line of the quasilinear portion of the recovery capacitive current $C_{m}\Phi$ $\approx$ $C_{m}\Phi_{H}$ intercepts the zero current axis at the point $V\approx V_{M}$ ) (See Fig. 9 and and Fig. 22). As we later apply the parsing of the rising edge ionic current into the activation mAvrami sodium current $J_{M}$ and the corresponding ionic polarization currents, our fits find that the reversal potential of the incoming activation current is $V_{M}$. The fact that reversal potentials of $J_{M}$, $J_{mM}$, $C_{m}\Phi_{M}$ and $C_{m}\Phi_{H}$ are the same supports the fact that the outgoing current $J_{H}$ is in fact sodium current. This result is also independent of any other assumption and it is solely determined by the charge conserving cable equation.
However, the current $J_{H}$ intercepts the zero current axis at $V=V_{H}$ that is significantly lower than the reversal potential $V_{M}$, and its polynomial trend line extension intercepts the slope of the recovery linear segment of the membrane current $J_{mH}$ at the potential $V_{M}$ but below the zero current axis. Fig. 10 displays the same behaviour at all temperatures. These facts support the conclusion that $J_{H}$ is in fact sodium current with its reversal potential modified from $V_{M}$ to $V_{H}$ by the inactivation currents and the continuous phase change of the sodium M-channel symmetry into the sodium H-channel symmetry while keeping almost all the mAvrami channels open (see Fig. 15).
Quasi linear segments of $J_{K}$ and $J_{N}$ and $J_{H}$, apparent at once in Figures 7, 8, 9, and SM Figures 21, 23, correspond to fractions of open channels equal to one; i.e. all channels open. The slopes of the capacitive quasi linear segments correspond to time rate parameters $\mu_{K}$, $\mu_{M}$, $\mu_{H}$, $\mu_{N}$, and the slopes of the ionic current quasi linear segments correspond to maximum conductance $g_{K}$, $g_{M}$, $g_{H}$ and $g_{N}$. Maximum conductance and time rates vs. Temperature are plotted in Fig. 12.
@@
Figure 11: The rising edge capacitive, membrane, ionic and corresponding polarizations, and mAvrami fitted currents are plotted against time. The three mAvrami fitted currents $C_{m}\Phi_{M}$, $J_{mM}$, $J_{M}$, and the $Mo(t)/M$ curve begin and intersect the zero-current axis at the inception point $t=to_{M}$, the time at which M-sodium channels start to open. Capacitive and membrane polarization currents $C_{m}\Phi_{MP}$ and $J_{mMP}$, start with the polarization flip at $t=to_{M}$ and then decay. The surface under the first segment of capacitive polarization curve $C_{m}\Phi_{MP}$ is equal to the polarization charge $Q_{P}\approx 10x10^{-9}$ Coulomb/cm2 moved across the membrane (a.k.a. gating charge). The flip from H-channel symmetry to M-channel symmetry symmetry and the transfer of charges across but within the membrane in conjunction with the ionic polarization current precedes the opening of sodium channels. In the present theory the motion of gating charges tapers off at approximately the maximum rate of rise of the action potential at about 60 mV when the fraction of open sodium channels is about $20\%$ . The value of the charge $Q_{P}$ remains approximately the same at all temperatures. Towards the peak of the action potential, preceding the flip from M-channel symmetry to H-channel symmetry at the peak, there is an opposite charge transfer across and also within the membrane amounting to $q_{P}\approx-2.6x10^{-9}$ Coulomb/cm2.
@@
Quasi linear segments of ionic currents $J_{K}$, $J_{M}$ and $J_{N}$ intersect the zero current axis at the same point as the corresponding quasi linear segments of capacitive and membrane currents. The cable Eq. (3) yields for these Quasi linear segments the relationships between maximum conductance and corresponding time rate. See also the reference [45]: